OpenStudy (anonymous):

Create two sets of three triangle lengths: one set that does create a right triangle and another that does not. Show and explain why the first set does create a right triangle and why the second set does not. I NEED HELP PLZZZ AND THANK U

OpenStudy (anonymous):

3,4,5 The sum f squares of shorter sides is equal to the square of the longest side. That is, \[3^{2}+4^{2}=5^{2}\] Therfore it forms a right triangle. 4,5,6 The sum f squares of shorter sides is equal to the square of the longest side. That is, \[4^{2}+5^{2}\neq6^{2}\] Therfore it dus nt form a right triangle.

OpenStudy (anonymous):

1)LHS=3^2+4^2 =9+16 =25 RHS=5^2 =25 LHS=RHS 2)LHS=4^2+5^2 =16+25 =41 RHS=6^2 =36 LHS is nt equal to RHS

OpenStudy (anonymous):

im confused which one is a right triangle and which isnt i have noidea what to do